OpenStudy (anonymous):

1. The compound RX3 decomposes according to the equation 3RX3  R + R2X3 + 3X2 In an experiment the following data were collected for the decomposition at 100°C. What is the average rate of reaction over the entire experiment? t(s) [RX3](mol L–1) 0 0.85 2 0.67 6 0.41 8 0.33 12 0.20 14 0.16

OpenStudy (aaronq):

the average of the rate is simply the summation of all the rates recorded over the number of measurements. \(\Delta rate=\dfrac{\Sigma ~ rate_i}{6}\)

OpenStudy (anonymous):

I'm sorry I still don't understand. Please simplify more.

OpenStudy (aaronq):

\(\sf \Delta rate =\dfrac{rate_1+rate_2...+rate_6}{number~of~measurements}\)

OpenStudy (anonymous):

I used your equation but it does not fit into any of the answer choices. A) 0.011 mol L–1s–1 B) 0.022 mol L–1s–1 C) 0.049 mol L–1s–1 D) 0.069 mol L–1s–1 E) 0.078 mol L–1s–1

OpenStudy (aaronq):

it wants it in terms of time, so after you found the average rate, divide by the time elapsed.

OpenStudy (anonymous):

ok I did this. .85+.67+.41+.33+.20+.16/6=0.44 then divided that by 14=0.031 is that correct?

OpenStudy (aaronq):

hmm thats what i think is right. i tried it like \(\Delta rate=\dfrac{\Delta rate}{\Delta t}=\dfrac{|rate_6-rate_1|}{t_6-t_1}=0.0492\) which is think is what they want, but it seems oversimplified to me.

OpenStudy (anonymous):

yeah i guess i misunderstood. But thank you that is very simplified.

OpenStudy (aaronq):

no, you were doing it fine, the answer they wanted is not what the actual average of the rate would be.

OpenStudy (aaronq):

you're welcome, nonetheless.